A solution was prepared by dissolving 0.1921 g of
Answer #1

HgO + 4Br- + H2O ------> HgBr4^-2 + 2OH-

Moles of HgO = 0.1921 /200.59 = 0.00096

Moles of KBr = 4 / 119 = 0.0336

1 mole of HgO reacts with 4 moles of Br- and forms 2 moles of OH-

0.00096 moles reacts with ==>0.00384 moles of Br- and forms 0.00192 moles of OH-

Moles of OH- = Moles of HCl

moles = molarity * volume

0.00192 = 17.96*M

M = 0.107 M